# Infinite-Width Limit of a Single Attention Layer: Analysis via Tensor Programs

### v1

@@ 1.1. Related Works
Two pioneering studies have investigated the NNGP of self-attention layers Yan19b ; HBSDN (20). Initially, Yan19b pointed out that an unconventional scaling factor of $1/n$ in the softmax function enables straightforward NNGP evaluation. Greg Yang has introduced the theoretical framework of Tensor Programs Yan19a ; Yan20a ; Yan20b , systematically composing kernel functions, including NNGPs, for modern neural networks. Transformers with the $1/n$-scaling fall within the scope of the applicability of Tensor Programs. However, the $1/n$-scaling is rarely used in theoretical or practical contexts, leaving the more realistic $1/\sqrt{n}$-scaling unresolved. To attack this problem, HBSDN (20) analyzed self-attention by varying the number of heads. They numerically demonstrated non-Gaussian behavior emerging in the single-head case due to stochasticity and correlations within the attention matrix, even at infinite embedding dimensions. They further showed that if we take the infinite limit of the number of heads, Gaussian behavior emerges in the self-attention output and defines an NNGP kernel termed infinite attention. Although this infinite attention empirically improved performance on certain NNGP regression benchmarks, its suitability as a theoretical foundation for realistic self-attention remains uncertain. This is because practical attention implementations typically use only 1-128 heads EXW+ (24), far fewer than the embedding dimension. In contrast to prior studies, our framework explicitly addresses scenarios with a finite number of heads and the resulting non-Gaussian effects.
2. Preliminary
2.1. Notation and Setup of Neural Networks
We define the notation for a standard neural network and its usage. In what follows, we denote by $n$ the dimensionality corresponding to the network’s width. Although we can vary $n$ across different layers or architectures, for simplicity we here treat every layer as having the same width $n$.
@@ 2.1. Notation and Setup of Neural Networks
Applying the row-wise SoftMax to the score matrix, we define
@@
We then weight the values to form each head’s output $\mathrm{Head}^{(a)}=A^{(a)}V^{(a)}\in\mathbb{R}^{s\times n}$. Finally, we sum across all $H$ heads to recover the full attention output $\mathrm{MultiHead}=\sum_{a=1}^{H}\mathrm{Head}^{(a)}(W^{O,a})^{\top}\in\mathbb{R}^{s\times n}$, where each row $\mathrm{MultiHead}_{i\cdot}$ is given by
$(\mathrm{MultiHead}_{i\cdot})^{\top}=\sum_{a=1}^{H}\sum_{j=1}^{s}W^{O,a}W^{V,a}x^{i}\mathrm{SoftMax}_{j}(p^{(a)}_{i,1},\dots,p^{(a)}_{i,s})\in\mathbb{R}^{n}\quad(i\in[s]).$ (4)
@@
This setup of weight matrices is common in practice (see GB (10) for summary) and also identical to that in the existing tensor networks Yan19a ; Yan19b ; Yan20b .
2.2. Distribution of Attention Mechanism in Previous Setup
@@
with some $\kappa>0$ (see Appendix A and Theorem E.8 in Yan19b for details). Intuitively, in the $1/n$-scaling regime, the dot-product score $p_{i,j}^{(a)}$ converges to 00 for every pair $(i,j)\in[s]^{2}$ and head $a\in[H]$, which simplifies the non-Gaussian behavior of the attention outputs.
@@
3. Main Theorem
3.1. Limiting Distribution
We introduce the limiting distribution for two types of quantities: variables generated within the Netsor program, and the scalar dot-products that arise in attention mechanisms. The convergence of these variables to their limiting distributions is formally established in our main result, Theorem 1.
Definition 1 (Limiting Distribution).
@@ 3.1. Limiting Distribution
    If $h$ is as an initial vector from $\mathcal{V}_{0}$, then $Z^{h}$ follows the distribution specified in Section 2.1.
@@
    $\mathrm{Cov}(Z^{g^{i}},Z^{g^{j}})=\begin{cases}0&(\text{ if }W^{i}\text{ and }W^{j}\text{ are different matrices}),\\ \sigma_{W}^{2}\mathbb{E}[Z^{h^{i}}Z^{h^{j}}]&(\text{ if }W^{i}\text{ and }W^{j}\text{ are the same matrix}).\end{cases}$
@@
In Definition 1, the formulations for the limiting random variables $Z^{h}$ associated with initial vectors, MatMul outputs, and Nonlin outputs are developed by the existing Tensor Programs framework (Yan19b, ). Specifically, the limiting distribution of initial vectors follows that given in Section 2.1; variables generated by MatMul converge to Gaussian distribution; and variables generated by Nonlin are nonlinear transformation of the corresponding limiting distributions.
@@
3.2. Statement
In this section, we show the convergence of the distribution of variables in the presence of an attention layer as the main result. Here, we refer to a function as pseudo-Lipschitz if it is pseudo-Lipschitz of order $d$ for some $d\in[2,\infty)$. See Definition 2 for the definition of pseudo-Lipschitz functions.
@@ 3.2. Statement
Fact 1 (Netsor Master Theorem Yan19b ).
@@
as $n\to\infty$. Here, $Z^{h^{1}},\dots,Z^{h^{k}}$ are defined in Definition 1, (i)–(iii).
@@
Theorem 1.
Consider a Netsor program, and suppose all nonlinearities used in Nonlin are pseudo-Lipschitz. We adopt the settings and notations from Assumption 1, which defines vectors $g^{1},\dots,g^{m}$ and scalar dot-products $p_{1},\dots,p_{r}$. Further, suppose all initial vectors and weight matrices are sampled as explained in Section 2.1. Now, let $h^{1},\dots,h^{k}\in\mathbb{R}^{n}$ be vectors whose elements are given by
@@
as $n\to\infty$. Here, $Z^{h^{j}}$ is given by $Z^{h^{j}}=\varphi^{j}(Z^{g^{1}},\dots,Z^{g^{m}},\mathring{p}_{1},\dots,\mathring{p}_{r})$ for $j\in[k]$, and $(Z^{g^{1}},\dots,Z^{g^{m}},\mathring{p}_{1},\dots,\mathring{p}_{r})$ is defined in Definition 1.
It may be noted that in the statement of Theorem 1, the boundedness of $\phi^{i,j}$ as in Assumption 1 is not essential, and we can drop this condition. However, we keep it for simplicity of the proof. Also, note that in this statement, $\mathring{p}_{1},\dots,\mathring{p}_{r}$ are random variables with positive variance, and the convergence in distribution ${\to}^{d}$ described in the statement refers to convergence to the distribution of them. Unlike the conventional master theorem derived from Tensor Programs Yan19a ; Yan19b , this theorem simultaneously characterizes the randomness of the ordinary variables like $g^{1},\dots,g^{m}$, along with the effects of finite-dimensional random variables $\mathring{p}_{1},\dots,\mathring{p}_{r}$.
Theorem 1 implies that, due to the randomness of $(\mathring{p}_{1},\dots,\mathring{p}_{r})$, the overall outputs resembles a hierarchical distribution, where the variance of one limiting distribution may depend on a realization of another limiting distribution $(\mathring{p}_{1},\dots,\mathring{p}_{r})$. This result differs from existing results by the Tensor Programs where the variance depends on moments of the other distributions. This result leads to a non-Gaussian limiting distribution in the attention case, whose details will be provided in Example 1.
@@
Corollary 2 (Coordinatewise Convergence).
Assume the same premise as in Theorem 1. Then, for all $\alpha\geq 1$, we have
@@
We present specific multi-head attention results as an example of the application of Corollary 2. This example implies that the output of the attention mechanisms is described by a non-Gaussian distribution. Specifically, when $y^{i}$ is an output of a multi-head attention layer as $\mathrm{MultiHead}_{i\cdot}$ defined in Eq. (4), $(Z^{y^{1}},\dots,Z^{y^{s}})$ follows a hierarchical Gaussian distribution whose conditioning $(\mathring{p}_{1},\dots,\mathring{p}_{r})=\{\mathring{p}_{i,j}^{(a)}:i,j\in[s],\ a\in[H]\}$ itself is a random variable, causing $(Z^{y^{1}},\dots,Z^{y^{s}})$ to converge to a heavy-tailed non-Gaussian distribution overall. We provide the details as follows:
Example 1 (Multi-Head Attention).
@@
$y_{\alpha}^{i}$ $=\sum_{a=1}^{H}\sum_{j=1}^{s}\mathrm{SoftMax}_{j}(p_{i,1}^{(a)},\dots,p_{i,s}^{(a)})\tilde{v}_{\alpha}^{a,j}$
@@
by introducing functions $\varphi^{i}\ (i\in[s])$, each of which is a pseudo-Lipschitz function. [^3] See Proposition 6 and Lemma 7. Note that $\mathrm{SoftMax}$ is Lipschitz so that it is also pseudo-Lipschitz (Fact 4). Then, by Corollary 2, we have
@@
where $Z^{y^{i}}$ is defined by $Z^{y^{i}}=\sum_{a=1}^{H}\sum_{j=1}^{s}\mathrm{SoftMax}_{j}(\mathring{p}_{i,1}^{(a)},\dots,\mathring{p}_{i,s}^{(a)})Z^{\tilde{v}^{a,j}}$. By Definition 1, $\{Z^{\tilde{v}^{a,j}}:j\in[s],\ a\in[H]\}$ and $\{\mathring{p}_{i,j}^{(a)}:i,j\in[s],\ a\in[H]\}$ satisfy the following.
@@
Note that $\{Z^{x^{j}}:j\in[s]\}$ can be computed by Definition 1.
4. Proof Sketch of Theorem 1
We sketch the proof of Theorem 1. First, observe that it is equivalent to the following statement:
Theorem 3.
Assume the same premise as in Theorem 1. Then, for any bounded and pseudo-Lipschitz function $\psi$, we have
$\frac{1}{n}\sum_{\alpha=1}^{n}\psi(g_{\alpha}^{1},\dots,g_{\alpha}^{m},p_{1},\dots,p_{r})\stackrel{{\scriptstyle d}}{{\longrightarrow}}\mathbb{E}[\psi(Z^{g^{1}},\dots,Z^{g^{m}},\mathring{p}_{1},\dots,\mathring{p}_{r})\mid\mathring{p}_{1},\dots,\mathring{p}_{r}]$
as $n\to\infty$, where $(Z^{g^{1}},\dots,Z^{g^{m}},\mathring{p}_{1},\dots,\mathring{p}_{r})$ is defined in Definition 1.
Clearly, Theorem 1 implies Theorem 3. Conversely, applying Proposition 6 establishes the opposite direction.
In what follows, we outline the proof of Theorem 3. Hereafter, $\bm{x}_{1:k}$ denotes the vector $(x_{1},\dots,x_{k})$, and likewise, $\bm{x}_{\alpha}^{1:k}$ denotes the vector $(x_{\alpha}^{1},\dots,x_{\alpha}^{k})$. Fix a bounded and pseudo-Lipschitz function $\psi$. By Lemma 5, it suffices to show that, for any bounded and Lipschitz function $f$,
@@ 4. Proof Sketch of Theorem 1
holds. Note that by Fact 1, we have
@@
for any pseudo-Lipschitz function $\tilde{\psi}$. Using this result, we first show that $\bm{p}_{1:r}$ converges in distribution to $\mathring{\bm{p}}_{1:r}$, which is a Gaussian vector defined by Definition 1 (Appendix C.1.1). Then, we consider
@@
We separately show $S_{1}\to 0$ (Appendix C.1.2) and $S_{2}\to 0$ (Appendix C.1.3).
5. Simulation and Discussion
We perform simulations to validate the infinite-width limit distributions derived in Example 1. For simplicity, the spatial dimension is fixed at $s=4$ in all experiments. More details on the simulation setting can be found in Appendix A. All simulation codes are available at https://github.com/manasakai/infinite-width-attention.
5.1. Effect of Finite Width

### v2

@@ 1.1 Related Works
Two pioneering studies have investigated the NNGP of self-attention layers Yan19b ; HBSDN (20). Initially, Yan19b pointed out that an unconventional scaling factor of $1/n$ in the softmax function enables straightforward NNGP evaluation. Greg Yang has introduced the theoretical framework of Tensor Programs Yan19a ; Yan20a ; Yan20b , systematically composing kernel functions, including NNGPs, for modern neural networks. Transformers with the $1/n$-scaling fall within the scope of the applicability of Tensor Programs. However, the $1/n$-scaling is rarely used in theoretical or practical contexts, leaving the more realistic $1/\sqrt{n}$-scaling unresolved. To attack this problem, HBSDN (20) analyzed self-attention by varying the number of heads. They numerically demonstrated non-Gaussian behavior emerging in the single-head case due to stochasticity and correlations within the attention matrix, even at infinite embedding dimensions. They further showed that if we take the infinite limit of the number of heads, Gaussian behavior emerges in the self-attention output and defines an NNGP kernel termed infinite attention. Although this infinite attention empirically improved performance on certain NNGP regression benchmarks, its suitability as a theoretical foundation for realistic self-attention remains uncertain. This is because practical attention implementations typically use only 1-128 heads EXW+ (24), far fewer than the embedding dimension.
Beyond the classical NNGP and Tensor Programs analyses, several recent works have examined Transformers under the standard $1/\sqrt{n}$ scaling and related asymptotic or dynamical regimes. DYZ (23) analyzed Transformers by tracking the first two moments (the kernel) of the signal to characterize propagation at initialization and during training. CNQG (24) applied mean-field theory to characterize the edge of chaos via forward and backward signal propagation, assuming Gaussianity of the QK product. BCP (24) employed dynamical mean field theory to study training dynamics under various infinite limits, including infinite width, heads, and depth, identifying parameterizations that ensure stable feature learning over time. NLL+ (23) modeled signal evolution with a stochastic differential equation, which requires modifying the softmax function.
2 Preliminary
2.1 Notation and Setup of Neural Networks
We define the notation for a standard neural network and its usage. In what follows, we denote by $n$ the dimensionality corresponding to the network’s width. Although we can vary $n$ across different layers or architectures, for simplicity we here treat every layer as having the same width $n$. A comprehensive summary of the notations used throughout this paper is provided in Appendix A.
@@ 2.1 Notation and Setup of Neural Networks
Here, the common scaling $1/\sqrt{n}$ in the definition of $p_{i,j}^{(a)}$ is a key issue in this study. [^1] Our analysis focuses on the $1/\sqrt{n}$-scaling, a choice motivated by its prevalence in both practical Transformer implementations and the theoretical literature (e.g., HBSDN (20); DYZ (23); CNQG (24)). While some analyses have raised questions about its stability YHB+ (21); BCP (24), the precise conditions for stable training remain an active area of research and may depend on factors often abstracted away in simplified limits, such as the number of tokens or data-specific statistics. For instance, recent work has shown that data statistics can significantly alter optimal scaling rules HL (25). Applying the row-wise SoftMax to the score matrix, we define
@@
We then weight the values to form each head’s output $\mathrm{Head}^{(a)}=A^{(a)}V^{(a)}\in\mathbb{R}^{s\times n}$. Finally, we sum across all $H$ heads to recover the full attention output $\mathrm{MultiHead}=H^{-\frac{1}{2}}\sum_{a=1}^{H}\mathrm{Head}^{(a)}(W^{O,a})^{\top}\in\mathbb{R}^{s\times n}$, where each row $\mathrm{MultiHead}_{i\cdot}$ is given by
$(\mathrm{MultiHead}_{i\cdot})^{\top}=\frac{1}{\sqrt{H}}\sum_{a=1}^{H}\sum_{j=1}^{s}W^{O,a}W^{V,a}x^{i}\mathrm{SoftMax}_{j}(p^{(a)}_{i,1},\dots,p^{(a)}_{i,s})\in\mathbb{R}^{n}\quad(i\in[s]).$ (4)
@@
This setup of weight matrices is common in practice (see GB (10) for summary) and also identical to that in the existing Tensor Programs Yan19a ; Yan19b ; Yan20b .
2.2 Distribution of Attention Mechanism in Previous Setup
@@
with some $\kappa>0$ (see Appendix A and Theorem E.8 in Yan19b for details). Intuitively, in the $1/n$-scaling regime, the dot-product score $p_{i,j}^{(a)}$ converges to 0 for every pair $(i,j)\in[s]^{2}$ and head $a\in[H]$, which simplifies the non-Gaussian behavior of the attention outputs.
@@
3 Main Theorem
3.1 Limiting Distribution
We introduce the limiting distribution for two types of quantities: variables generated within the Netsor program, and the scalar dot-products that arise in attention mechanisms. The convergence of these variables to their limiting distributions is formally established in our main result, Theorem 3.1.
Definition 3.1 (Limiting Distribution).
@@ 3.1 Limiting Distribution
    If $h$ is an initial vector from $\mathcal{V}_{0}$, then $Z^{h}$ follows the distribution specified in Section 2.1.
@@
[…]
@@
In Definition 3.1, the formulations for the limiting random variables $Z^{h}$ associated with initial vectors, MatMul outputs, and Nonlin outputs are developed by the existing Tensor Programs framework (Yan19b, ). Specifically, the limiting distribution of initial vectors follows that given in Section 2.1; variables generated by MatMul converge to Gaussian distribution; and variables generated by Nonlin are nonlinear transformations of the corresponding limiting distributions.
@@
3.2 Statement
In this section, we show the convergence of the distribution of variables in the presence of an attention layer as the main result. Here, we refer to a function as pseudo-Lipschitz if it is pseudo-Lipschitz of order $d$ for some $d\in[2,\infty)$. See Definition C.1 for the definition of pseudo-Lipschitz functions.
@@ 3.2 Statement
Fact 3.1 (Netsor Master Theorem Yan19b ).
@@
as $n\to\infty$. Here, $Z^{h^{1}},\dots,Z^{h^{k}}$ are defined in Definition 3.1, (i)–(iii).
@@
Theorem 3.1.
Consider a Netsor program, and suppose all nonlinearities used in Nonlin are pseudo-Lipschitz. We adopt the settings and notations from Assumption 3.1, which defines vectors $g^{1},\dots,g^{m}$ and scalar dot-products $p_{1},\dots,p_{r}$. Further, suppose all initial vectors and weight matrices are sampled as explained in Section 2.1. Now, let $h^{1},\dots,h^{k}\in\mathbb{R}^{n}$ be vectors whose elements are given by
@@
as $n\to\infty$. Here, $Z^{h^{j}}$ is given by $Z^{h^{j}}=\varphi^{j}(Z^{g^{1}},\dots,Z^{g^{m}},\mathring{p}_{1},\dots,\mathring{p}_{r})$ for $j\in[k]$, and $(Z^{g^{1}},\dots,Z^{g^{m}},\mathring{p}_{1},\dots,\mathring{p}_{r})$ is defined in Definition 3.1.
It may be noted that in the statement of Theorem 3.1, the boundedness of $\phi^{i,j}$ as in Assumption 3.1 is not essential, and we can drop this condition. However, we keep it for simplicity of the proof. Also, note that in this statement, $\mathring{p}_{1},\dots,\mathring{p}_{r}$ are random variables with positive variance, and the convergence in distribution ${\to}^{d}$ described in the statement refers to convergence to the distribution of them. Unlike the conventional master theorem derived from Tensor Programs Yan19a ; Yan19b , this theorem simultaneously characterizes the randomness of the ordinary variables like $g^{1},\dots,g^{m}$, along with the effects of finite-dimensional random variables $\mathring{p}_{1},\dots,\mathring{p}_{r}$.
Theorem 3.1 implies that, due to the randomness of $(\mathring{p}_{1},\dots,\mathring{p}_{r})$, the overall outputs resemble a hierarchical distribution, where the variance of one limiting distribution may depend on a realization of another limiting distribution $(\mathring{p}_{1},\dots,\mathring{p}_{r})$. This result differs from existing results by the Tensor Programs where the variance depends on moments of the other distributions. This result leads to a non-Gaussian limiting distribution in the attention case, whose details will be provided in Example 3.1.
@@
Corollary 3.2 (Coordinatewise Convergence).
Assume the same premise as in Theorem 3.1. Then, for all $\alpha\geq 1$, we have
@@
We present specific multi-head attention results as an example of the application of Corollary 3.2. This example implies that the output of the attention mechanisms is described by a non-Gaussian distribution. Specifically, when $y^{i}$ is an output of a multi-head attention layer as $\mathrm{MultiHead}_{i\cdot}$ defined in Eq. (4), $(Z^{y^{1}},\dots,Z^{y^{s}})$ follows a hierarchical Gaussian distribution whose conditioning $(\mathring{p}_{1},\dots,\mathring{p}_{r})=\{\mathring{p}_{i,j}^{(a)}:i,j\in[s],\ a\in[H]\}$ itself is a random variable, causing $(Z^{y^{1}},\dots,Z^{y^{s}})$ to converge to a heavy-tailed non-Gaussian distribution overall. We provide the details as follows:
Example 3.1 (Multi-Head Attention).
@@
$y_{\alpha}^{i}$ $=\frac{1}{\sqrt{H}}\sum_{a=1}^{H}\sum_{j=1}^{s}\mathrm{SoftMax}_{j}(p_{i,1}^{(a)},\dots,p_{i,s}^{(a)})\tilde{v}_{\alpha}^{a,j}$
@@
by introducing functions $\varphi^{i}\ (i\in[s])$, each of which is a pseudo-Lipschitz function. [^4] See Proposition C.3 and Lemma C.4. Note that $\mathrm{SoftMax}$ is Lipschitz so that it is also pseudo-Lipschitz (Fact C.3). Then, by Corollary 3.2, we have
@@
where $Z^{y^{i}}$ is defined by $Z^{y^{i}}=H^{-\frac{1}{2}}\sum_{a=1}^{H}\sum_{j=1}^{s}\mathrm{SoftMax}_{j}(\mathring{p}_{i,1}^{(a)},\dots,\mathring{p}_{i,s}^{(a)})Z^{\tilde{v}^{a,j}}$. By Definition 3.1, $\{Z^{\tilde{v}^{a,j}}:j\in[s],\ a\in[H]\}$ and $\{\mathring{p}_{i,j}^{(a)}:i,j\in[s],\ a\in[H]\}$ satisfy the following.
@@
Note that $\{Z^{x^{j}}:j\in[s]\}$ can be computed by Definition 3.1.
The pseudocode for this example is presented in Algorithm 1 in Appendix B.1.
Remark 3.1.
It is well-known in probability theory that hierarchical mixtures of Gaussian components typically lead to distributions with heavier tails than standard Gaussians. This observation highlights a significant challenge for theoretical analysis, since many existing frameworks, such as NNGP and Tensor Programs, are developed under regimes where the limiting distributions are Gaussian (or can be expressed as Gaussian components with correction terms). Extending these tools to rigorously handle such heavy-tailed non-Gaussian behaviors remains an important open problem, and we believe our work provides a concrete starting point for such future developments.
Remark 3.2.
Our result in Example 3.1 has a close relationship with the result of HBSDN (20). First, the limiting distribution of the attention scores in our framework perfectly matches the Gaussian distribution in their infinite-head limit. Second, the variance of our finite-head non-Gaussian output distribution theoretically matches the variance of their infinite-head Gaussian output distribution. This alignment suggests that while the infinite-head limit correctly captures the second-order statistics of the output, our finite-head analysis provides a more complete picture by characterizing the non-Gaussian aspects of the distribution that are not captured by second-order statistics alone.
4 Proof Sketch of Theorem 3.1
We sketch the proof of Theorem 3.1. First, observe that it is equivalent to the following statement:
Theorem 4.1.
Assume the same premise as in Theorem 3.1. Then, for any bounded and pseudo-Lipschitz function $\psi$, we have
$p$
as $n\to\infty$, where $(Z^{g^{1}},\dots,Z^{g^{m}},\mathring{p}_{1},\dots,\mathring{p}_{r})$ is defined in Definition 3.1.
Clearly, Theorem 3.1 implies Theorem 4.1. Conversely, applying Proposition C.3 establishes the opposite direction.
In what follows, we outline the proof of Theorem 4.1. Hereafter, $\bm{x}_{1:k}$ denotes the vector $(x_{1},\dots,x_{k})$, and likewise, $\bm{x}_{\alpha}^{1:k}$ denotes the vector $(x_{\alpha}^{1},\dots,x_{\alpha}^{k})$. Fix a bounded and pseudo-Lipschitz function $\psi$. By Lemma C.2, it suffices to show that, for any bounded and Lipschitz function $f$,
@@
holds. Note that by Fact 3.1, we have
@@
for any pseudo-Lipschitz function $\tilde{\psi}$. Using this result, we first show that $\bm{p}_{1:r}$ converges in distribution to $\mathring{\bm{p}}_{1:r}$, which is a Gaussian vector defined by Definition 3.1 (Appendix D.1.1). Then, we consider
@@
We separately show $S_{1}\to 0$ (Appendix D.1.2) and $S_{2}\to 0$ (Appendix D.1.3).
5 Simulation and Discussion
We perform simulations to validate the infinite-width limit distributions derived in Example 3.1. Details on the simulation setting can be found in Appendix B. All simulation codes are available at https://github.com/manasakai/infinite-width-attention.
5.1 Effect of Finite Width